I am a 2-digit number. The sum of my digits is 6.
My units digit is 2 times my tens digit. Who am I?

Answer by Edwin McCravy(20081)

I am a 2-digit number.
Let t = your tens digit
Let u = your units digit

The sum of my digits is 6.
    t + u = 6

My units digit is 2 times my tens digit.
        u = 2t

So you have this system of equations:

    t + u = 6
        u = 2t

Substitute 2t for u in t + u = 6

   t + 2t = 6

       3t = 6
     
        t = 2

u = 2t = 2(2) = 4

Who am I?
You are the number 24. 

Edwin